Hello
I'm stumped on this question. How do I do about solving it?
Sandy and Aleesha have some money
3/4 of Sandra's money is equal to 2/5 of Aleesha's money
Aleesha has $490 more than Sandy
How much money does Sandy have
Thanks in advance.

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

3/4 of Sandra's money it equal to 2/5 of Aleesha's money
Aleesha has $490 more than Sandy

let S = sandra's money.
let A = aleesha's money.

3/4 * S = 2/5 * A
A = S + 490

replace A with S + 490 in the first equation to get"
3/4 * S = 2/5 * (S + 490
simplify to get:
3/4 * S = 2/5 * S + 196
subtract 2/5 * S from both sides of the equation to get:
3/4 * S - 2/5 * S = 196
3/4 * 5/5 = 15/20
2/5 * 4/4 = 8/20
equation becomes:
15/20 * S - 8/20 * S = 196
combine like terms to get:
7/20 * S = 196
solve for S to get:
S = 20/7 * 196 = 560.

replace S with 560 in the equation of 3/4 * S = 2/5 * A to get:
3/4 * 560 = 2/5 * A
multiply both sides of the equation by 5/2 to get:
5/2 * 3/4 * 560 = A
solve for A to get:
A = 1050.

you have S = 560 and A = 1050.
3/4 * S = 2/5 * A becomes:
3/4 * 560 = 2/5 * 1050 which becomes:
420 = 420.
this confirms the solution is correct.

the solution is sandy has 560 dollars.

Answer by Alan3354(69443)

Sandy and Aleesha have some money
3/4 of Sandra's money is equal to 2/5 of Aleesha's money
Aleesha has $490 more than Sandy
How much money does Sandy have
------------------
S*3/4 = A*2/5
15S = 8A
---
A = S + 490
=================
15S = 8A
15S = 8*(S + 490) = 8S + 3920
7S = 3920
S = $560